# RoVerFly: Robust and Versatile Learning-based Control of Quadrotor Across Payload Configurations

### v1

@@ Abstract
Designing robust controllers for precise, arbitrary trajectory tracking with quadrotors is challenging due to nonlinear dynamics and underactuation, and becomes harder with flexible cable–suspended payloads that introduce extra degrees of freedom and hybridness. Classical model-based methods offer stability guarantees but require extensive tuning and often do not adapt when the configuration changes—when a payload is added or removed, or when the payload mass or cable length varies. We present RoVerFly, a unified learning-based control framework in which a reinforcement learning (RL) policy serves as a robust and versatile tracking controller for standard quadrotors and for cable-suspended payload systems across a range of configurations. Trained with task and domain randomization, the controller is resilient to disturbances and varying dynamics. It achieves strong zero-shot generalization across payload settings—including no payload as well as varying mass and cable length—without controller switching or re-tuning, while retaining the interpretability and structure of a feedback tracking controller. Code and supplementary materials are available at https://https://github.com/mintaeshkim/roverfly.
@@ I Introduction
Quadrotors are widely used for aerial navigation, and numerous planning and control strategies have been developed for agile and precise maneuvering [1, 2]. For payload transport, rigid (active) attachment increases manipulation capability but reduces agility due to higher inertia [3], whereas cable suspension preserves maneuverability and enables multi-UAV cooperation at the cost of additional underactuation and nonlinear dynamics [4, 5].
@@
Prior work on quadrotor–payload systems falls into three directions: (1) generating and tracking payload trajectories that suppress swing [6, 7, 8]; (2) controlling systems with elastic/flexible cables and pulley mechanisms [9, 10, 11]; and (3) cooperative manipulation with multiple quadrotors [7, 10]. Most approaches use model-based geometric control leveraging differential flatness [4, 12]. While these methods offer theoretical guarantees, they require accurate models and extensive tuning, and often do not generalize to dynamic or uncertain settings. A key limitation is the need for separate controller designs for different configurations (with vs. without payload) [1, 13], which necessitates controller switching in practice and can introduce instability. Moreover, reliance on flatness requires precomputed trajectories and precludes tracking dynamically infeasible references.
@@
Model-free RL has emerged as a promising paradigm for UAV control, learning feedback policies directly from interaction without explicit dynamics modeling [14, 2, 15, 16]. This is particularly compelling for UAVs, whose high-rate, nonlinear, underactuated, and aerodynamically coupled dynamics are difficult to model and tune accurately [17, 18]. RL controllers have shown robust trajectory tracking [19], cross-platform adaptation [20], disturbance rejection [21], and the ability to track dynamically infeasible trajectories [15]. However, neither existing RL nor model-based methods adequately handle rapidly varying payload conditions such as changes in mass or cable length, and there is no RL controller explicitly designed for flexible cable-suspended payload tracking.
@@
Table I summarizes the landscape. Racing-oriented learning controllers excel on fixed courses with specialized perception/estimation stacks but are not intended for arbitrary references or payload handling [2]. General-purpose RL benchmarking suites enable reproducible studies yet encourage task-specific rewards and typically train a separate policy per task rather than a single versatile controller [22]. Hybrid learning–plus–adaptive control can follow references and reject wind but relies on a nominal model and an additional adaptive layer, with demonstrations mainly on relatively simple, planar motions [15]. Methods that adapt via latent extrinsics capture cross-platform variation near hover but require an adaptation phase and are evaluated chiefly under small-impulse disturbances [20]. Taken together, these limitations point to the need for a controller that tracks arbitrary references while accommodating unknown payload presence, cable length, and mass—without per-task policies or controller switching.
@@
We introduce a learning-based tracking controller that operates across varying payload and cable configurations—using a single policy with no switching or re-tuning. Task and domain randomization make the policy robust to underactuation, nonlinear dynamics, changing task conditions, and disturbances. Drawing on hybrid dynamics and POMDP insights, we use I/O history as a belief-state proxy so the policy acts as an implicit observer under partial observability. The result is a single RL-learned feedback law that unifies hovering, takeoff, landing, and arbitrary trajectory tracking while avoiding task-specific reward engineering or environment-specific tuning.
@@
Contributions. (i) Unified control framework: a single policy for robust arbitrary trajectory tracking in both quadrotor-only and flexible cable-suspended payload systems, generalizing across payload mass and cable length without fine-tuning. (ii) RL for flexible payload tracking: an versatile RL-based controller explicitly targeting flexible cable-suspended payloads with strong zero-shot performance via task and domain randomization. (iii) Belief-based use of I/O history: a justification for incorporating I/O history grounded in POMDP belief propagation and hybrid dynamics analysis, showing its role in reducing uncertainty and improving performance under partial observability.
@@
TABLE I: Comparison of learning-based quadrotor control approaches.
@@ II-A Model-Based Control and Trajectory Planning for Quadrotor-Payload Systems
Model-based approaches—particularly geometric control—provide formulations with almost-global stability guarantees for quadrotor–payload systems [4, 6, 7, 12]. Via differential flatness, they plan trajectories by expressing states and inputs as flat outputs, ensuring feasibility. Geometric controllers on this principle track both quadrotor and payload trajectories [4, 1]. Recent work extends these methods to elastic or flexible cables [9, 10], introducing higher-order dynamics and greater underactuation. However, they require accurate models, frequent retuning, and are sensitive to unmodeled effects (e.g., actuator saturation, external disturbances), limiting robustness in practice.
Trajectory-generation methods likewise plan dynamically feasible payload motions. Optimization-based approaches such as mixed-integer quadratic programming (MIQP) [6] and mathematical programs with complementarity constraints (MPCC) [8] handle hybrid dynamics and cable-tension regimes for swing suppression and obstacle avoidance. MIQP is computationally heavy, while MPCC improves efficiency but remains model- and tuning-sensitive; as a result, these methods are typically used offline or in settings with limited real-time adaptability.
@@ II-B Learning-based Control of Quadrotors
Model-free RL has become a practical tool for quadrotor flight, enabling high-performance trajectory tracking and agile maneuvers without explicit system identification. In this view, the learned policy acts as a nonlinear feedback controller that adapts to complex dynamics from data. RL controllers show disturbance rejection when paired with adaptive elements [15], improved sample efficiency and sim-to-real transfer with recent training strategies [23], and race-level performance where simulation-trained controllers surpassed human pilots [2].
@@
Despite this progress, payload trajectory tracking—especially with flexible, cable-suspended payloads—remains underexplored. One exception learns a dynamics model from human demonstrations and embeds it in MPC for stabilization [24], but targets near-static, vertically suspended loads, limiting dynamic maneuvering. Other high-performance results rely on precise estimation stacks and specialized hardware yet are restricted to fixed racing courses, limiting generalization [2]. Simulation frameworks offer task flexibility via modular rewards but typically require training a separate policy per task [22].
@@
Robustness to hardware/parameter variation has been addressed by learning latent extrinsics such as mass or thrust [20], though these approaches require an adaptation phase and are mostly validated near hover. Hybrid RL with adaptive control improves robustness and can track arbitrary references, but still assumes partially known dynamics and is shown in constrained scenarios [15]. These gaps motivate a unified, general-purpose learning-based controller that serves as a single robust tracker across diverse quadrotor tasks—including payload manipulation—without task-specific design, controller switching, or retraining.
@@
III Problem Description
III-A System Dynamics
We consider two quadrotor variants: (1) a standard platform, and (2) a quadrotor with a flexible cable-suspended payload where cable length and payload mass may vary. Although their configurations and observations differ, both admit the same control structure and can be handled within a unified learning-based framework.
Our primary model is the quadrotor–payload system, which reduces to the quadrotor-only case as $m_{P}\to 0$ and the cable constraint becomes inactive. With a taut, massless cable, the system has eight degrees of freedom, configuration space $SE(3)\times S^{2}$, and four degrees of underactuation. The dynamics are:
$\mathbf{x}_{Q}$ $=\mathbf{x}_{P}-l\,\mathbf{q},$ (1)
$\dot{\mathbf{x}}_{P}$ $=\mathbf{v}_{P},$ (2)
$\quad+\mathbf{w}_{F},$ (3)
$\dot{\mathbf{q}}$ $=\boldsymbol{\omega}\times\mathbf{q},$ (4)
$\dot{\mathbf{R}}$ $=\mathbf{R}\,\hat{\boldsymbol{\Omega}}$, (6)
This formulation captures the coupled translation–rotation behavior and the nonlinear coupling between thrust direction and payload swing. Removing the payload yields a standard $SE(3)$ quadrotor model; the same RL-learned feedback law applies with a reduced observation space. Using this generalized model, we aim to learn a single policy that robustly tracks reference trajectories across configurations without redesigning model-specific controllers.
III-B Arbitrary Trajectory Tracking
We pose a robust reference–tracking problem under dynamic and uncertain conditions. The reference $\mathbf{x}_{\mathrm{ref}}(t)$ is sampled from a family of sinusoid-sum trajectories with randomized amplitudes, frequencies, and phases, and includes hovering intervals at the start and end.
Trained on this distribution, the controller shows zero-shot generalization to unseen references from the same family. The task contains sharp turns and dynamically infeasible segments, where flatness-based geometric controllers typically fail due to the need for precomputed flat outputs.
Robustness is evaluated via:
  • External disturbances: impulse forces in $0.5,0.5]~\mathrm{N}$ and torques in $0.005,0.005]~\mathrm{N\cdot m}$ applied for $0,0.5]~\mathrm{s}$ during hover.
  • Initial-state perturbations: random offsets in position, orientation, and linear/angular velocity for both quadrotor and payload.
Across cases, the controller rejects disturbances, stabilizes to hover, and resumes accurate tracking, serving as a robust feedback tracker for both system configurations.
IV-A Preliminaries
We formulate the problem as a POMDP $\mathcal{M}$, defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\gamma,r)$, where an agent interacts with an environment with state space $\mathcal{S}$ and action space $\mathcal{A}$, governed by the transition function $\mathcal{T}(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$. The agent receives an observation $\mathbf{o}_{t}\in\mathcal{O}$ through the observation function $O(\mathbf{o}_{t}|\mathbf{s}_{t},\mathbf{a}_{t-1})$, which depends on the current state and the previous action. Episodes start from an initial-state distribution $\rho_{0}$ over $\mathcal{S}$, i.e., $\mathbf{s}_{0}\sim\rho_{0}$. Given a reward function $r(\mathbf{o}_{t},\mathbf{a}_{t})$ and a discount factor $\gamma$, the objective in reinforcement learning is to maximize the expected discounted return: $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\mathbf{o}_{t},\mathbf{a}_{t})\right]$. The goal is to learn an optimal policy $\pi^{*}\in\Pi$ that maps observations $\mathbf{o}_{t}\in\mathcal{O}$ to optimal actions $\mathbf{a}_{t}^{*}\in\mathcal{A}$.
IV-B Algorithm and Training Overview
We train a policy $\pi_{\theta}$ that maps observations to actions at $100\,\text{Hz}$. The observation comprises present, past, and future parts: present features $(\mathbf{s}_{t},\mathbf{e}_{t},\mathbf{a}_{t-1})$; a short I/O history $\boldsymbol{\zeta}_{t}$ of recent states, references, and actions over the last $H$ steps; and a feedforward term $\boldsymbol{\xi}_{t}$ with previewed position/velocity in the body frame. As in Fig. 1, the history and feedforward stacks pass through optional MLP encoders to form embeddings, which are concatenated with the present features and fed to $\pi_{\theta}$. The actor and critic are separate two-layer MLPs ($\times256$) with SiLU activations. The policy is trained with PPO in a MuJoCo environment [25, 26].
Figure 1: Block diagram of the policy and training loop. Present features are concatenated with history/preview embeddings and fed to $\pi$. CTBR commands pass through the rate loop and mixer before actuation in MuJoCo.
IV-C State, Action, and Observation Spaces
IV-C1 State

### v2

# RoVerFly: Robust and Versatile Implicit Hybrid Control of Quadrotor-Payload Systems

@@ Abstract
Designing robust controllers for precise trajectory tracking with quadrotors is challenging due to nonlinear dynamics and underactuation, and becomes harder with flexible cable–suspended payloads that add degrees of freedom and hybrid dynamics. Classical model-based methods offer stability guarantees but require extensive tuning and often fail to adapt when the configuration changes—when a payload is added or removed, or when its mass or cable length varies. We present RoVerFly, a unified learning-based control framework where a single reinforcement learning (RL) policy functions as an implicit hybrid controller, managing complex dynamics without explicit mode detection or controller switching. Trained with task and domain randomization, the controller is resilient to disturbances and varying dynamics. It achieves strong zero-shot generalization across payload settings—including no payload as well as varying mass and cable length—without re-tuning, while retaining the interpretability and structure of a feedback tracking controller. Code and supplementary materials are available at https://github.com/mintaeshkim/roverfly.
@@ I Introduction
Quadrotors are widely used for aerial navigation, and numerous planning and control strategies have been developed for agile, precise maneuvering [1, 2]. For payload transport, rigid (active) attachment increases manipulation capability but reduces agility due to higher inertia [3], whereas cable suspension preserves maneuverability and enables multi-UAV cooperation at the cost of additional underactuation, nonlinear dynamics, and hybrid dynamics from cable slack/taut transitions [4, 5].
@@
Prior work on quadrotor–payload systems falls into three directions: (1) generating and tracking payload trajectories that suppress swing [6, 7, 8]; (2) controlling systems with elastic/flexible cables and pulley mechanisms [9, 10, 11]; and (3) cooperative manipulation with multiple quadrotors [7, 10]. Most approaches use model-based geometric control leveraging differential flatness [4, 12]. While these methods offer theoretical guarantees, they require accurate models and extensive tuning, and often fail to generalize to dynamic or uncertain settings. A key limitation is the need for separate controller designs for different configurations (with vs. without payload) [1, 13], which necessitates explicit switching in practice and can introduce instability. Moreover, reliance on flatness requires precomputed trajectories and precludes tracking dynamically infeasible references.
@@
Model-free RL has emerged as a promising paradigm for UAV control, learning feedback policies directly from interaction without explicit dynamics modeling [14, 2, 15, 16]. This is particularly compelling for UAVs, whose high-rate, nonlinear, underactuated, and aerodynamically coupled dynamics are difficult to model and tune accurately [17, 18]. RL controllers have shown robust trajectory tracking [19], cross-platform adaptation [20], disturbance rejection [21], and tracking of dynamically infeasible trajectories [15]. However, existing methods adequately handle rapidly varying payload conditions such as changes in mass or cable length, nor do they offer a unified solution to the switching problem inherent in these hybrid systems.
@@
Table I summarizes the landscape. Racing-oriented learning controllers excel on fixed courses with specialized perception/estimation stacks but are not intended for arbitrary references or payload handling [2]. General-purpose RL benchmarking suites enable reproducible studies yet encourage task-specific rewards and typically train a separate policy per task rather than a single versatile controller [22]. Hybrid learning–plus–adaptive control can follow references and reject wind but relies on a nominal model and an additional adaptive layer, with demonstrations mainly on simple planar motions [15]. Methods that adapt via latent extrinsics capture cross-platform variation near hover but require an adaptation phase and are evaluated chiefly under small disturbances [20]. Taken together, these limitations point to the need for a controller that tracks arbitrary references while accommodating unknown payload presence, cable length, and mass—without per-task policies or switching.
@@
We introduce RoVerFly, a learning-based tracking controller whose core is a single policy functioning as an implicit hybrid controller. This data-driven approach directly addresses the key limitation of classical methods by learning to manage the varied, hybrid dynamics of the payload system—from no payload to a flexible suspended cable—without explicit mode detection or switching. By leveraging task and domain randomization, the policy becomes robust to underactuation, nonlinearities, and disturbances. Drawing on POMDP insights, we use I/O history as a belief-state proxy, enabling the policy to act as an implicit observer to infer latent system properties and make robust decisions under partial observability. The result is a single RL-learned feedback law that unifies hovering, takeoff, landing, and trajectory tracking while avoiding task-specific reward engineering or environment-specific tuning.
@@
Contributions. (i) Unified, implicit hybrid control framework: A single policy functioning as an implicit hybrid controller for robust arbitrary trajectory tracking in both quadrotor-only and flexible cable-suspended payload systems, generalizing across payload mass and cable length without switching or fine-tuning. (ii) Data-driven control for flexible cable payloads: A versatile learning-based controller targeting the complex underactuated dynamics of flexible cable-suspended payloads, with strong zero-shot performance achieved via task and domain randomization. (iii) Belief-based control using I/O history: A justification for incorporating I/O history grounded in POMDPs, showing its role in enabling the policy to implicitly estimate the system’s hybrid mode and state for robust control under partial observability.
@@
TABLE I: Comparison of existing learning-based quadrotor control approaches.
@@ II-A Model-Based Control and Trajectory Planning for Quadrotor-Payload Systems
Model-based approaches—particularly geometric control—provide formulations with almost-global stability guarantees for quadrotor–payload systems [4, 6, 7, 12]. Via differential flatness, they plan trajectories by expressing states and inputs as flat outputs, ensuring feasibility. Geometric controllers on this principle track both quadrotor and payload trajectories [4, 1]. Recent work extends these methods to elastic or flexible cables [9, 10], introducing higher-order dynamics and added underactuation. However, they require accurate models, frequent retuning, and are sensitive to unmodeled effects (e.g., actuator saturation, disturbances), limiting robustness in practice.
Trajectory-generation methods likewise plan dynamically feasible payload motions. Optimization-based approaches such as mixed-integer quadratic programming (MIQP) [6] and mathematical programs with complementarity constraints (MPCC) [8] handle hybrid dynamics and cable-tension regimes for swing suppression and obstacle avoidance. MIQP is computationally heavy, while MPCC improves efficiency but remains model- and tuning-sensitive; as a result, these methods are typically used offline or in settings with limited adaptability.
@@ II-B Learning-based Control of Quadrotors
Model-free RL has become a practical tool for quadrotor flight, enabling high-performance trajectory tracking and agile maneuvers without explicit system identification. The learned policy acts as a nonlinear feedback controller adapting to complex dynamics from data. RL controllers show disturbance rejection when paired with adaptive elements [15], improved sample efficiency and sim-to-real transfer with recent strategies [23], and race-level performance where simulation-trained controllers surpassed human pilots [2].
@@
Despite progress, payload trajectory tracking—especially with flexible, cable-suspended payloads—remains underexplored. One exception learns a dynamics model from human demonstrations and embeds it in MPC for stabilization [24], but targets near-static, vertically suspended loads, limiting dynamic maneuvering. Other high-performance results rely on precise estimation stacks and specialized hardware restricted to fixed racing courses, limiting generalization [2]. Simulation frameworks offer task flexibility via modular rewards but typically require training a separate policy per task [22].
@@
Robustness to hardware/parameter variation has been addressed by learning latent extrinsics such as mass or thrust [20], though these require an adaptation phase and are mostly validated near hover. Hybrid RL with adaptive control improves robustness and can track arbitrary references, but still assumes partially known dynamics and is shown in constrained scenarios [15]. These gaps motivate a unified, general-purpose learning-based controller serving as a single robust tracker across diverse quadrotor tasks—including payload manipulation—without task-specific design, switching, or retraining.
Figure 1: Diagram of the training loop. Present features are concatenated with history/preview embeddings and fed to $\pi$. Collective thrust and body rates (CTBR) commands pass through the rate loop and mixer before actuation.
III Problem Description
III-A System Dynamics
We consider two quadrotor variants: (1) a standard platform, and (2) a quadrotor with a flexible cable-suspended payload where cable length and payload mass may vary. Although configurations and observations differ, both admit the same control structure and can be handled within a unified learning-based framework.
The dynamics of the quadrotor-payload system are hybrid, with two operational modes: taut and slack. The active mode depends on the cable tension $T$, determined by the system state.
III-A1 Taut Mode Dynamics ($T>0$)
When the cable is taut, the payload is constrained and the dynamics are tightly coupled. For a taut, massless cable, the system has eight degrees of freedom, configuration space $SE(3)\times S^{2}$, and four degrees of underactuation. The dynamics are [4]:
$\mathbf{x}_{Q}$ $=\mathbf{x}_{P}-l\,\mathbf{q},$ (1)
$\dot{\mathbf{x}}_{P}$ $=\mathbf{v}_{P},$ (2)
$\quad+\mathbf{w}_{F},$ (3)
$\dot{\mathbf{q}}$ $=\boldsymbol{\omega}\times\mathbf{q},$ (4)
$\dot{\mathbf{R}}$ $=\mathbf{R}\,\hat{\boldsymbol{\Omega}}$, (6)
III-A2 Slack Mode Dynamics and Switching Conditions
When the cable is slack ($T=0$, i.e., $\|\mathbf{x}_{Q}-\mathbf{x}_{P}\|<l$), quadrotor and payload dynamics decouple. The quadrotor behaves as a standard $SE(3)$ system governed by
$m_{Q}\ddot{\mathbf{x}}_{Q}=f\mathbf{R}\mathbf{e}_{3}-m_{Q}g\mathbf{e}_{3}+\mathbf{w}_{F},$ (8)
while the payload acts as a free-falling point mass,
$m_{P}\ddot{\mathbf{x}}_{P}=-m_{P}g\mathbf{e}_{3}.$ (9)
The rotational dynamics remain as in Eq. (6)-(7).
Mode transitions follow state-dependent guard conditions. Slack→taut occurs when the cable becomes taut ($\|\mathbf{x}_{Q}-\mathbf{x}_{P}\|\geq l$), representing a non-smooth impact. Taut→slack occurs when tension $T$ drops to zero.
This explicit hybrid structure, with non-smooth transitions, challenges traditional model-based control, which typically requires supervisory logic to detect the active mode and switch between laws. In contrast, our work learns a single, continuous policy that acts as an implicit hybrid controller. It does not monitor switching conditions but learns a robust strategy from interaction data that navigates transitions. While the above analysis uses an idealized model for clarity, all simulations are conducted in MuJoCo with a finite-stiffness cable, ensuring the learned policy is validated on a realistic flexible model capturing slack–taut dynamics.
III-B Arbitrary Trajectory Tracking
We pose a robust reference–tracking problem under dynamic, uncertain conditions. The reference $\mathbf{x}_{\mathrm{ref}}(t)$ is sampled from a family of sinusoid trajectories with randomized amplitudes, frequencies, and phases, and with hovering intervals at the start and end. Trained on this distribution, the controller shows zero-shot generalization to unseen references from the same family. The task contains sharp turns and dynamically infeasible segments, where flatness-based controllers typically fail due to precomputed flat outputs. Robustness is evaluated under two conditions: external disturbances, modeled as impulse forces in $0.5,0.5]~\mathrm{N}$ and torques in $0.005,0.005]~\mathrm{N\cdot m}$ applied for $0,0.5]~\mathrm{s}$; and initial-state perturbations, introduced as random offsets in position, orientation, and linear/angular velocity for both quadrotor and payload. Across cases, the controller rejects disturbances, stabilizes hover, and resumes accurate tracking, serving as a robust feedback tracker for both system configurations.
IV-A Preliminaries
We formulate the problem as a POMDP $\mathcal{M}$, defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\gamma,r)$, where an agent interacts with an environment with state space $\mathcal{S}$ and action space $\mathcal{A}$, governed by transition function $\mathcal{T}(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$. The agent receives observation $\mathbf{o}_{t}\in\mathcal{O}$ through the observation function $O(\mathbf{o}_{t}|\mathbf{s}_{t},\mathbf{a}_{t-1})$, which depends on the current state and previous action. Episodes start from an initial distribution $\rho_{0}$ over $\mathcal{S}$, i.e., $\mathbf{s}_{0}\sim\rho_{0}$. Given reward $r(\mathbf{o}_{t},\mathbf{a}_{t})$ and discount factor $\gamma$, the objective in RL is to maximize the expected discounted return: $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\mathbf{o}_{t},\mathbf{a}_{t})\right]$. The goal is to learn an optimal policy $\pi^{*}\in\Pi$ mapping observations $\mathbf{o}_{t}\in\mathcal{O}$ to optimal actions $\mathbf{a}_{t}^{*}\in\mathcal{A}$.
IV-B Algorithm and Training Overview
We train a policy $\pi_{\theta}$ mapping observations to actions at $100\,\text{Hz}$. The observation includes present, past, and future parts: present features $(\mathbf{s}_{t},\mathbf{e}_{t},\mathbf{a}_{t-1})$; a short I/O history $\boldsymbol{\zeta}_{t}$ of recent states, references, and actions over $H$ steps; and a feedforward term $\boldsymbol{\xi}_{t}$ with previewed position/velocity in the body frame. As in Fig. 1, the history and feedforward stacks pass through MLP encoders to form embeddings, concatenated with present features and fed to $\pi_{\theta}$. The actor and critic are separate two-layer MLPs ($\times256$) with SiLU activations. The policy is trained with PPO [25].
@@ IV-B Algorithm and Training Overview
IV-C Belief-Based Analysis of I/O History for Implicit Hybrid Control
The hybrid nature of the quadrotor–payload system presents a significant challenge in partial observability. The discrete operational mode—taut or slack—is a latent variable that is not directly measured and dictates the system’s underlying dynamics. Inferring this mode and the full continuous state $\mathbf{s}_{t}$ from a single observation $\mathbf{o}_{t}$ is intractable. To address this, our policy uses a fixed-length history of observations and actions (I/O history). This does not violate the Markov property but rather restores it by providing a practical surrogate for the true belief state over both the discrete mode and continuous state. This history allows the policy to function as an implicit observer, learning to infer the system’s hybrid state and enabling the stable, continuous control actions characteristic of an implicit hybrid controller.
IV-C1 Belief-Based Policy and Value under Uncertainty
This update reduces entropy by exploiting temporal context:
Lower-entropy beliefs improve performance under:
Hence, belief accuracy is critical in POMDPs.
IV-C2 Hybrid Dynamics and Implicit Mode Estimation
The core challenge for our I/O history-based approach is estimating the unobserved hybrid mode $m_{t}$. For the flexible cable system, dynamics can be modeled as:
where mode $m_{t}$ is unobserved. An accurate belief over the mode is critical, as the optimal action can differ drastically between taut and slack regimes. The agent must maintain a joint belief over mode and state:
Updating this belief requires marginalizing over past modes and states:
$\qquad\times b_{t-1}(m_{t-1},\mathbf{s}_{t-1})\,d\mathbf{s}_{t-1}.$
Because mode transitions depend on latent variables (like cable tension), a single observation $\mathbf{o}_{t}$ is insufficient; temporal context is needed to disambiguate modes. Entropy shrinks with longer histories:
$>H(m_{t},\mathbf{s}_{t}\mid\mathbf{o}_{0:t},\mathbf{a}_{0:t-1}).$
This grounding shows why I/O history is the key mechanism enabling our policy to act as an implicit hybrid controller. By providing temporal context, the history lets the policy form a richer internal representation of the system’s belief state, reducing uncertainty over both discrete mode ($m_{t}$) and continuous state ($\mathbf{s}_{t}$). This ability to implicitly infer hybrid dynamics without an explicit observer or switching logic allows a single, continuous policy to achieve robust control across non-smooth transitions. The importance of this mechanism is empirically validated by our ablation studies in Table IV.
IV-D State, Action, and Observation Spaces
IV-D1 State